A combustion device was used to analyze a 0.6349 g compound containing only carbon, hydrogen, and oxygen. The experiment produced 1.603 g carbon dioxide and 0.2810 g of water. Determine the empirical formula of the compound


1
Expert's answer
2021-01-11T03:50:25-0500

0.281g H2O ("\frac{2g}{18g}") = 0.0312gH

and 1.603gCO2 ("\frac{12g}{44g}") = 0.437gC

0.6349gTotal - 0.437gC - 0.0312gH = 0.167gO

0.437gC("\frac{1mol}{12g}") = 0.0364molC "\implies" "\frac{0.0364molC}{0.0104}" = 3.5 (2) = 7C

0.0312gH("\frac{1mol}{1g}") = 0.0312molH "\implies" "\frac{0.0312molH}{0.0104}" = 3 (2) = 6H

0.167gO("\frac{1mol}{16g}") = 0.0104molO "\implies" "\frac{0.0104molO}{0.0104}" = 1(2) = 2O

the empirical formula of the compound is C7H6O2